Mike is in the back of a truck moving due east at a constant velocity of 20 m/s relative to the ground, if Mike throws an apple straight out the back of the truck (west) at a velocity of 15 m/s relative to the truck,

what is the velocity of the apple relative to the ground?

   m/s

What is the velocity of the apple relative to the ground if Mike throws the apple towards the front of the truck (East)?

   m/s
Answer #1

A]  velocity of the apple relative to the ground = 20 - 15 = 5 m/s due East

B] velocity of the apple relative to the ground = 20 + 15 = 35 m/s due East